The length of the base of an isosceles triangle is one third the length of one of its legs. If the perimeter of the triangle is 56 inches, what is the length of the base.
Answer by rfer(16322)

2 1/3x=56
7/3x=56
x=56/(7/3)
x=56*3/7
x=168/7
x=24 in legs
base =8 in